A 1500-kg vehicle travels at a constant speed of 22 m/s around a circular track that has a radius of 85 m.

[A] What is the magnitude of the acceleration of the vehicle?

[B] What is the average velocity of the vehicle during one revolution?

[C] What is the period (T) of the vehicle?

[D] What is the magnitude of the centripetal force?
Answer #1

Part A.

Centripetal acceleration is given by:

ac = V^2/R

V = 22 m/sec

R = 85 m

So,

ac = 22^2/85

ac = 5.69 m/sec^2

Part B.

Average velocity is given by:

Avg velocity = displacement/time taken

Since in one revolution, displacement = 0

So,

Avg Velocity = 0 m/sec

Part C.

Period of vehicle is given by:

T = 2*pi/w

w = angular velocity = V/R

So,

T = 2*pi*R/V

T = 2*pi*85/22

T = 24.3 sec

Part D.

Centripetal force will be

Fc = m*ac

Fc = 1500*5.69

Fc = 8535 N

Please Upvote.

Comment below if you've any query.